I have to complete the square for this equation:
(x-7)^2-100=0
I think you have to add 100 to both sides but I don't know what do to from there. I have to get the x
Please help soon!

Answer by Alan3354(69443)

I have to complete the square for this equation:
(x-7)^2-100=0
I think you have to add 100 to both sides but I don't know what do to from there. I have to get the x
-------------------
Add 100
(x-7)^2 = 10^2
Take the sqrt of both sides
x-7 = 10, and
x-7 = -10
-------------
x = 17
x = -3